How many tons of hay at $125 per ton and $75 per ton must be purchased to have 100 tons of hay worth $9000?
Found 3 solutions by checkley75, stanbon, Kalmetam:
Answer by checkley75(3666)

125x+75(100-x)=9000
125x+7500-75x=9000
50x=9000-7500
50x=1500
x=1500/50
x=30 tons of $125 hay.
100-30=70 tons of $75 hay.
proof
125*30+75*70=9000
3750+5250=9000
9000=9000

Answer by stanbon(75887)

How many tons of hay at $125 per ton and $75 per ton must be purchased to have 100 tons of hay worth $9000?
----------------------
Let # of tons of $125 hay be "x" ; value of that is 125x dollars.
------------------------
# of tons of $75 hay is "100-x" ; value of that is 75(100-x)=7500-75x dollars
-------------------------
EQUATION:
value + value = 9000 dollars
125x + 7500-75x = 9000
50x = 1500
x = 30 (# of tons of $125 per ton hay that must be purchased)
100-x = 70 (# of tons of $75 per ton hay that must be purchased)
================
Cheers,
Stan H.

Answer by Kalmetam(43)

Ok this is not a logic question so lets try it out...
125x72=9000 and 72 tons
75x120=9000 and 120 tons

That is the basic, we're not down to the answer.. yet
we need a little educated guess, so lets try 16 tons of $125 and 84 tons of $75.

125x16=2000
75x84=6300
2000+6300=8300
Pretty close.. the difference between 125 and 75 is 50 and we need 700 more so:
700/50=14
That means we need to turn 14 tons of $75 into $125 tons.
So that means we should have 30 tons of $125 and 70 tons of of $75
125x30=3750
75x70=5250
3750+5250=9000
30+70=100
Oh we now found the answer.. there were 30 tons worth $125 ans 70 tons worth $75.